# Quantum error correction near the coding theoretical bound

### v3

@@ 2 Introduction
Protograph codes [8] are LDPC codes whose parity-check matrices are constructed using a protograph matrix composed of permutation matrices (PM) as submatrices. A subclass of these codes includes quasi-cyclic (QC) LDPC codes [9], which are constructed using circulant permutation matrices (CPMs), and APM-LDPC (affine permutation matrices) codes [10], which are constructed using APMs. Similar to unstructublack LDPC codes, protograph codes are also known for their excellent error correction performance.
Short cycles are particularly known to cause a high error floor. The error floor of LDPC codes refers to a phenomenon where the error rate does not decrease as rapidly as expected with decreasing noise levels, especially in the high-fidelity regime. In computational error correction, extremely low error rates are requiblack, making it critically important to address the error floor phenomenon. It is known that the girth (the length of the shortest cycle) of QC-LDPC matrices is upper-bounded by 12 [9]. However, APM-LDPC matrices can exceed this upper bound on girth [10, 11].
LDPC codes are typically defined over a binary field. Certain types of LDPC codes are known to achieve further improvements in error-correcting performance by applying extensions to non-binary finite fields [12], although this increases the computational complexity requiblack for decoding.
@@
CSS codes defined by orthogonal sparse parity-check matrix pairs are referblack to as LDPC-CSS codes. Since LDPC-CSS codes are decoded using algorithms similar to those for classical LDPC codes, they are particularly promising for practical application in large-scale quantum computers. Algebraic construction methods allow for the straightforward construction of orthogonal pairs of QC-LDPC matrices [15].
@@
To date, in previous studies, no quantum error-correcting codes approaching the hashing bound and decodable with computational complexity proportional to the number of physical qubits had been discoveblack. This study aims to construct LDPC-CSS codes that simultaneously achieve a error floor down to FER $10^{-4}$, approach the hashing bound, and are decodable with computational complexity proportional to the number of physical qubits.
@@
The remainder of this paper is organized as follows. In Section 3.1, we describe the proposed methods for code construction and decoding. Section 3.5 presents the error correction performance results of the proposed codes. Further details of these sections are provided in the following section.
@@ 3.2 Permutation Matrices
For $J=2$, $L=4$, and $P=12$, we present an example of the above construction method. The function arrays $\underline{f}=(5x+4,\ 5x+8)$ and $\underline{g}=(7x+6,\ 7x+9)$, constructed using APMs, satisfy the requiblack conditions. In this case, the girth of the resulting parity-check matrices $\hat{H}_{X}$ and $\hat{H}_{Z}$ is 8. Figure 1 shows the matrices $\hat{H}_{X}$ and $\hat{H}_{Z}$ derived from $\underline{f}$ and $\underline{g}$, where black cells indicate ones and white cells indicate zeros.
@@ 3.4 Decoding method
In a depolarizing channel, the probabilities of X, Y, and Z errors are each given by $p_{D}/3$. In this case, the marginal probability that either X or Z error occurs is $f_{m}=2p_{D}/3$. The decoding result is defined as successful if the decoding algorithm correctly estimates both error vectors: $\underline{x}=\hat{\underline{x}},\underline{z}=\hat{\underline{z}}$. If the estimation fails for even a single bit of $\underline{x}$ or $\underline{z}$, the decoding is consideblack failed. The frame error rate (FER) is defined as the ratio of the number of experiments where at least one bit was incorrectly estimated to the total number of experiments.
Note that we do not take degenerate errors [21] into consideration; they are treated as decoding failures. Properly addressing degeneracy is generally consideblack essential for approaching the Hashing bound. To close the remaining gap to the Hashing limit, it will be necessary to explicitly handle degeneracy. However, a rigorous treatment of degeneracy is beyond the scope of this work and is left for future study.
@@ 3.5 Numerical Results
The CSS codes defined by $(H_{X},H_{Z})$ are of coding rate R=12J/LR=1-2J/L. As a result, the coding rates of the CSS codes constructed above are $R\in\{0.50,0.60,0.75\}$. For the CSS codes defined by $(H_{X},H_{Z})$, the quantum error-correction performance over a depolarizing channel was evaluated. The number of physical qubit is given by $n=ePL$. The parameters of the codes used in the experiments are listed in Table 1. Here, $n$ denotes the number of physical qubits, $k$ is the number of logical bits, and $\overline{d}$ represents an upper bound on the minimum distance. Since determining the exact minimum distance is computationally difficult, we estimated an upper bound using the following approach.
Table 1: Parameters of the proposed codes • and the code $\circ$ [16] that was previously consideblack to be the closest to the hashing bound in [16].
@@
In Figure 2, the decoding performance $(f_{m},\mathrm{FER})$ is plotted for each rate $R$ (from left to right, $R=0.75,0.60,0.50$) by varying the parameter $P$, where the number of physical qubits $n=ePL$ is proportional to $P$. Furthermore, high error-floor was not observed, at least down to an FER of $10^{-4}$. No conventional quantum LDPC codes have been reported where the FER shows error floors down to FER $10^{-4}$ and waterfall near the hashing bound [22, 23]. We also compare our results with those of Kasai et al. (2011), which were obtained under similar parameters. As indicated by the gray arrow, a significant performance improvement is observed.
Although this paper focuses on the case $L\geq 8$, recent results [24] have shown that relatively high error floors can emerge at large block lengths when $L=6$. To address this issue, our recent work [24] demonstrates that degeneracy-aware post-processing, which specifically targets stopping sets, can significantly blackuce the FER in the low-noise regime.
@@
Figure 3 plots $(f_{m}^{*},R)$ where $f_{m}^{*}$ denotes the value of $f_{m}$ to achieve FER of $10^{-4}$, and $R$ is the quantum coding rate. It can be observed that all rates approach the hashing bound. There is no conventional quantum error correction that achieves both an error floor down to FER $10^{-4}$ and steep waterfall near the hashing bound [23]. The results show that the proposed codes exhibit decoding performance that approaches the hashing bound.
Figure 2: Decoding performance of the proposed code: $f_{m}$ vs. FER. Decoding performance $(f_{m},\mathrm{FER})$ of the proposed code with $e=8$, $J=2$, $L\in\{8,10,16\}$, $P\in\{32,128,1024,8192\}$, and $R\in\{0.50,0.60,0.75\}$. The black solid line indicates the hashing bound. The code length, that is, the number of physical bits, is given by $n=ePL$. It is evident that the proposed method simultaneously achieves both error floors down to FER $10^{-4}$ and waterfall near the hashing bound. The gray dashed curve represents a rate-1/2 code [16] that was previously consideblack to be the closest to the hashing bound [22].
Figure 3: Decoding performance of the proposed code: $f_{m}$ vs. $R$. The plot shows the pairs $(f_{m}^{*},R)$, where $f_{m}^{*}$ denotes the value of $f_{m}$ to achieve FER of $10^{-4}$, and $R$ is the quantum coding rate. It can be observed that all rates approach the hashing bound. The filled circles represent the results of the proposed method, while the empty circles correspond to the method defined over the same finite field as proposed in [16].
@@ 5.1 Finite field extension to $(H_{\Gamma},H_{\Delta})$
The matrix in equation (10) can be transformed into its blackuced row echelon form using elementary row operations without requiring division. By randomly selecting a random solution of this system, we obtain $\underline{\lambda}$ followed by $\underline{\gamma}$ and $H_{\Gamma}$. Furthermore, treating the non-zero entries of $H_{\Delta}$ as variables, the linear equation $H_{\Gamma}H_{\Delta}^{\mathsf{T}}=0$ is solved to determine the non-zero entries of $H_{\Delta}$.
@@ 5.4 Details of decoding methods
then $(\hat{x}_{j})$ and $(\hat{z}_{j})$ are accepted as the estimated noise, and the algorithm terminates. If the syndromes are not satisfied, a decoding failure is declablack.

### v4

@@ 2 Introduction
Protograph codes [8] are LDPC codes whose parity-check matrices are constructed using a protograph matrix composed of permutation matrices (PM) as submatrices. A subclass of these codes includes quasi-cyclic (QC) LDPC codes [9], which are constructed using circulant permutation matrices (CPMs), and APM-LDPC (affine permutation matrices) codes [10], which are constructed using APMs. Similar to unstructured LDPC codes, protograph codes are also known for their excellent error correction performance.
Short cycles are particularly known to cause a high error floor. The error floor of LDPC codes refers to a phenomenon where the error rate does not decrease as rapidly as expected with decreasing noise levels, especially in the high-fidelity regime. In computational error correction, extremely low error rates are required, making it critically important to address the error floor phenomenon. It is known that the girth (the length of the shortest cycle) of QC-LDPC matrices is upper-bounded by 12 [9]. However, APM-LDPC matrices can exceed this upper bound on girth [10, 11].
LDPC codes are typically defined over a binary field. Certain types of LDPC codes are known to achieve further improvements in error-correcting performance by applying extensions to non-binary finite fields [12], although this increases the computational complexity required for decoding.
@@
CSS codes defined by orthogonal sparse parity-check matrix pairs are referred to as LDPC-CSS codes. Since LDPC-CSS codes are decoded using algorithms similar to those for classical LDPC codes, they are particularly promising for practical application in large-scale quantum computers. Algebraic construction methods allow for the straightforward construction of orthogonal pairs of QC-LDPC matrices [15].
@@
To date, in previous studies, no quantum error-correcting codes approaching the hashing bound and decodable with computational complexity proportional to the number of physical qubits had been discovered. This study aims to construct LDPC-CSS codes that simultaneously achieve a error floor down to FER $10^{-4}$, approach the hashing bound, and are decodable with computational complexity proportional to the number of physical qubits.
@@
The remainder of this paper is organized as follows. In “Results”, we describe the proposed methods for code construction and decoding. “Numerical Results” presents the error correction performance results of the proposed codes. Further details of these sections are provided in the following section.
@@ 3.2 Permutation matrices
For $J=2$, $L=4$, and $P=12$, we present an example of the above construction method. The function arrays $\underline{f}=(5x+4,\ 5x+8)$ and $\underline{g}=(7x+6,\ 7x+9)$, constructed using APMs, satisfy the required conditions. In this case, the girth of the resulting parity-check matrices $\hat{H}_{X}$ and $\hat{H}_{Z}$ is 8. Figure 1 shows the matrices $\hat{H}_{X}$ and $\hat{H}_{Z}$ derived from $\underline{f}$ and $\underline{g}$, where black cells indicate ones and white cells indicate zeros.
@@ 3.4 Decoding method
In a depolarizing channel, the probabilities of X, Y, and Z errors are each given by $p_{D}/3$. The decoding result is defined as successful if the decoding algorithm correctly estimates both error vectors: $\underline{x}=\hat{\underline{x}},\underline{z}=\hat{\underline{z}}$. If the estimation fails for even a single bit of $\underline{x}$ or $\underline{z}$, the decoding is considered failed. The frame error rate (FER) is defined as the ratio of the number of experiments where at least one bit was incorrectly estimated to the total number of experiments.
Note that we do not take degenerate errors [21] into consideration; they are treated as decoding failures. Properly addressing degeneracy is generally considered essential for approaching the Hashing bound. To close the remaining gap to the Hashing limit, it will be necessary to explicitly handle degeneracy. However, a rigorous treatment of degeneracy is beyond the scope of this work and is left for future study.
@@ 3.5 Numerical results
The CSS codes defined by $(H_{X},H_{Z})$ are of coding rate R=12J/LR=1-2J/L. As a result, the coding rates of the CSS codes constructed above are $R\in\{0.50,0.60,0.75\}$. For the CSS codes defined by $(H_{X},H_{Z})$, the quantum error-correction performance over a depolarizing channel was evaluated. The number of physical qubit is given by $n=ePL$. The parameters of the codes used in the experiments are listed in Table 1. Here, $n$ denotes the number of physical qubits, $k$ is the number of logical bits, and $\overline{d}$ represents an upper bound on the minimum distance. Since determining the exact minimum distance is computationally difficult, we estimated an upper bound using the following approach. We enumerated all the shortest cycles in the matrices and used the minimum weight of the codewords derived from these cycles as the upper bound on the minimum distance.
Table 1: Parameters of the proposed codes • and the code $\circ$ [16] that was previously considered to be the closest to the hashing bound in [16].
@@
In Figure 2, the decoding performance $(p_{D},\mathrm{FER})$ is plotted for each rate $R$ (from left to right, $R=0.75,0.60,0.50$) by varying the parameter $P$, where the number of physical qubits $n=ePL$ is proportional to $P$. Furthermore, high error-floor was not observed, at least down to an FER of $10^{-4}$. No conventional quantum LDPC codes have been reported where the FER shows error floors down to FER $10^{-4}$ and waterfall near the hashing bound [22, 23]. We also compare our results with those of Kasai et al. (2011), which were obtained under similar parameters. As indicated by the gray arrow, a significant performance improvement is observed.
Although this paper focuses on the case $L\geq 8$, recent results [24] have shown that relatively high error floors can emerge at large block lengths when $L=6$. To address this issue, our recent work [24] demonstrates that degeneracy-aware post-processing, which specifically targets stopping sets, can significantly reduce the FER in the low-noise regime.
@@
Figure 3 plots $(p_{D}^{*},R)$ where $p_{D}^{*}$ denotes the value of $p_{D}$ to achieve FER of $10^{-4}$, and $R$ is the quantum coding rate. It can be observed that all rates approach the hashing bound. There is no conventional quantum error correction that achieves both an error floor down to FER $10^{-4}$ and steep waterfall near the hashing bound [23]. The results show that the proposed codes exhibit decoding performance that approaches the hashing bound.
Figure 2: Decoding performance of the proposed code: $p_{D}$ vs. FER. Decoding performance $(p_{D},\mathrm{FER})$ of the proposed code with $e=8$, $J=2$, $L\in\{8,10,16\}$, $P\in\{32,128,1024,8192\}$, and $R\in\{0.50,0.60,0.75\}$. The black solid line indicates the hashing bound. The code length, that is, the number of physical bits, is given by $n=ePL$. It is evident that the proposed method simultaneously achieves both error floors down to FER $10^{-4}$ and waterfall near the hashing bound. The gray dashed curve represents a rate-1/2 code [16] that was previously considered to be the closest to the hashing bound [22].
Figure 3: Decoding performance of the proposed code: $p_{D}$ vs. $R$. The plot shows the pairs $(p_{D}^{*},R)$, where $p_{D}^{*}$ denotes the value of $p_{D}$ to achieve FER of $10^{-4}$, and $R$ is the quantum coding rate. It can be observed that all rates approach the hashing bound. The filled circles represent the results of the proposed method, while the empty circles correspond to the method defined over the same finite field as proposed in [16].
@@ 5.1 Finite field extension to $(H_{\Gamma},H_{\Delta})$
The matrix in equation (10) can be transformed into its reduced row echelon form using elementary row operations without requiring division. By randomly selecting a random solution of this system, we obtain $\underline{\lambda}$ followed by $\underline{\gamma}$ and $H_{\Gamma}$. Furthermore, treating the non-zero entries of $H_{\Delta}$ as variables, the linear equation $H_{\Gamma}H_{\Delta}^{\mathsf{T}}=0$ is solved to determine the non-zero entries of $H_{\Delta}$.
@@ 5.3 Properties of the companion matrix
The binary representations of the vectors $\underline{v}(\alpha^{i})$ and $\underline{w}(\alpha^{i})$, along with the corresponding companion matrices $A^{i}$ and $(A^{\mathsf{T}})^{i}$ for a primitive element $\alpha\in\mathbb{F}_{8}$, are summarized in Fig. 4.
@@ 5.4 Details of decoding methods
then $(\hat{x}_{j})$ and $(\hat{z}_{j})$ are accepted as the estimated noise, and the algorithm terminates. If the syndromes are not satisfied, a decoding failure is declared.